# Overcoming Uncertain Incompleteness for Robust Multimodal Sequential Diagnosis Prediction via Curriculum Data Erasing Guided Distillation

### v3

@@ 1 Introduction
Missing data, a common issue in reality, exacerbates model performance. Basic approaches, such as imputation by mean or excluding incomplete data instances, often fail to preserve true data distribution and result in information loss [5]. Advanced statistical techniques, including Multivariate Imputation by Chained Equations (MICE) [6], show better efficacy, but their application in complex multimodal scenarios still remains challenging. Therefore, deep learning approaches such as reconstruction [7, 8, 9], which impute missing features, and knowledge distillation (KD) [10], which transfers teacher’s knowledge on complete data to a student learning with incomplete data [11, 12], have gained popularity.
@@ 2.2.2 Systematic Knowledge Distillation Framework
$\mathcal{L}_{\text{MWCD}}=\sum_{m\{D,N,C\}}\{\alpha\mathcal{L}_{\text{MWCD}}^{\text{\text{teacher}, $\rightarrow$ student},m}+(1-\alpha)\mathcal{L}_{\text{MWCD}}^{\text{student $\rightarrow$ \text{teacher}, },m}\}$ (3)
@@ 2.2.3 Curriculum Learning Guided Random Data Erasing
Firstly, the teacher is trained using curriculum learning guided random data erasing, starting with easier samples and gradually progressing to more difficult ones. Each modality is assigned a missing probability of 0.0 or 0.1 with equal probability until specific epoch $e_{1}$, after which the probability of 0.2 is added. Thereafter, during the distillation, complete data representations from the teacher trained in the previous manner are migrated until epoch $e_{2}$, after which training continues with either no missing data or a 0.1 missing ratio.
@@ 3.1 Experimental Setup
To handle missing data, we assign a value beyond the existing range in demographics and diagnosis codes. For instance, if the total number of codes is 3882, the missing value is assigned as 3883. We also replace missing tokens in clinical notes with the UNK token [9].

### v4

# Overcoming Uncertain Incompleteness for Robust Multimodal Sequential Diagnosis Prediction via Curriculum Data Erasing Guided Knowledge Distillation

@@ 1 Introduction
Missing data, a common issue in reality, exacerbates model performance [5]. Basic approaches, such as imputation by mean or excluding incomplete data instances, often fail to preserve true data distribution and result in information loss [6]. Advanced statistical techniques, including Multivariate Imputation by Chained Equations (MICE) [7], show better efficacy, but their application in complex multimodal scenarios still remains challenging. Therefore, deep learning approaches such as reconstruction [8, 9, 10], which impute missing features, and knowledge distillation (KD) [11], which transfers teacher’s knowledge on complete data to a student learning with incomplete data [12, 13], have gained popularity.
@@ 2.2.2 Systematic Knowledge Distillation Framework
$\mathcal{L}_{\text{MWCD}}=\sum_{m\in\{D,N,C\}}\{\alpha\mathcal{L}_{\text{MWCD}}^{\text{\text{teacher}, $\rightarrow$ student},m}+(1-\alpha)\mathcal{L}_{\text{MWCD}}^{\text{student $\rightarrow$ \text{teacher}, },m}\}$ (3)
@@ 2.2.3 Curriculum Learning Guided Random Data Erasing
Firstly, the teacher is trained using curriculum learning guided random data erasing, starting with easier samples and gradually progressing to more difficult ones. All modalities are assigned a missing probability of 0.0 or 0.1 with equal probability until specific epoch $e_{1}$, after which the probability of 0.2 is added. Thereafter, during the distillation, complete data representations from the teacher trained in the previous manner are migrated until epoch $e_{2}$, after which training continues with either no missing data or a 0.1 missing ratio to each modality.
@@ 3.1 Experimental Setup
To handle missing data, we assign a value beyond the existing range in demographics and diagnosis codes. For instance, if the total number of codes is 3882, the missing value is assigned as 3883. We also replace missing tokens in clinical notes with UNK token [10].